OpenStudy (anonymous):

36.5% =36.5/100 =365/1000 =73/200 =0.365 how do they get 73 and 200????please help!1

OpenStudy (anonymous):

percentage to fraction or decimal....

OpenStudy (phi):

divide 365 by 5 what do you get ?

OpenStudy (anonymous):

thanks Phi!!

OpenStudy (anonymous):

i get 73!!

OpenStudy (phi):

in other words, if you can divide the top and the bottom by the same number, you can simplify (reduce) the fraction \[ \frac{365 }{ 1000} =\frac{73\cdot 5 }{ 200 \cdot 5} =\frac{73 }{ 200 } \cdot \frac{5}{5}= \frac{73 }{ 200 } \cdot 1= \frac{73 }{ 200 } \]

OpenStudy (anonymous):

aww i get now! thanks again...really appreciate the help :)

OpenStudy (phi):

yw